What is the mole fraction of sodium chloride and water in a solution, prepared by mixing 256 g of sodium chloride and 176 g water? (Na=23g/mol Cl= 35g/mol H=1g/mol o=16g/mol)


1
Expert's answer
2022-05-11T10:28:03-0400

"M(NaCl)=23+35=58 g\/mol"

"n(NaCl)=256\/58=4.4 mol"

"M(H_2O)=2(1)+16=18 g\/mol"

"n(H_2O)=176\/18=9.8 mol"

"w(NaCl)=n(NaCl)\/(n(NaCl)+n(H_2O))=4.4\/(4.4+9.8)=0.31"

"w(H_2O)=n(H_2O)\/(n(NaCl)+n(H_2O))=9.8\/(4.4+9.8)=0.69"